Family Tree

Family Tree

About Me

My photo
Kathmandu, Bagmati Zone, Nepal
I am Basan Shrestha from Kathmandu, Nepal. I use the term 'BASAN' as 'Balancing Actions for Sustainable Agriculture and Natural Resources'. I am a Design, Monitoring & Evaluation professional. I hold 1) MSc in Regional and Rural Development Planning, Asian Institute of Technology, Thailand, 2002; 2) MSc in Statistics, Tribhuvan University (TU), Kathmandu, Nepal, 1995; and 3) MA in Sociology, TU, 1997. I have more than 10 years of professional experience in socio-economic research, monitoring and documentation on agricultural and natural resource management. I had worked in Lumle Agricultural Research Centre, western Nepal from Nov. 1997 to Dec. 2000; CARE Nepal, mid-western Nepal from Mar. 2003 to June 2006 and WTLCP in far-western Nepal from June 2006 to Jan. 2011, Training Institute for Technical Instruction (TITI) from July to Sep 2011, UN Women Nepal from Sep to Dec 2011 and Mercy Corps Nepal from 24 Jan 2012 to 14 August 2016 and CAMRIS International in Nepal commencing 1 February 2017. I have published articles to my credit.

Thursday, June 13, 2019

Mutually Exclusive and Inclusive Events: Examples, Statistical Note 41

Mutually Exclusive Events: Two events are mutually exclusive if they cannot occur at the same time. In tossing of a coin, either of head or tail flips meaning they are mutually exclusive. An occurrence of head excludes an occurrence of tail in the same toss. 








In another example of drawing a card from the deck of 52 cards, the occurrence of an ace card excludes cards of other ranks, two to ten, jack, queen and king.








Mutually Inclusive Events: Two events are mutually inclusive or non-exclusive if the occur of an event does not exclude the occurrence of another event at the same time. In drawing a card from the deck of 52 cards, the occurrence of an ace card does not exclude the occurrence of a club card. Because, both identities - ace as a rank and club as a suit exist on the same card.

Saturday, May 25, 2019

Heads and Runs in Multiple Tossing of a Coin and Contingency Table, Statistical Note 40

List the possible outcomes in three tosses of a coin, categorize them by the number of number of heads and the number of runs in the sequence and prepare a contingency table.

Tossing of a coin thrice has two to the power three equal to eight possible outcomes (refer to my statistical note 9). Every outcome is measured using two discrete random variables – the number of heads and the number of runs. The outcomes have number of heads ranging from zero to three as mutually exclusive categories of response as shown in Table 1. A run is a sequence of flips of the same face of a coin. The number of runs ranges from one to three as mutually exclusive categories of response (Table 1). Example, an outcome HTH has three runs, as every toss has a different face than the previous toss. A contingency table, also known as the cross tabulation, crosstab or two-way table counts the number observations for each category of two variables.

Table 1: Possible outcomes in three tosses of a coin by number of heads and number of runs










The contingency table presents the number of possible outcomes by number of heads and number of runs (Table 2).







Central values of both variables have high probability of occurrence. Example, outcomes with one and two heads respectively have higher probability of occurrence, three by eight. Likewise, outcomes with two runs have higher probability of occurrence, four by eight. Refer to my statistical note 3 for probability and contingency table.

Tuesday, May 14, 2019

Multiplication Rule of Dependent Events, Statistical Note 39

If two cards are drawn without replacement from a deck of cards, what is the joint probability that both cards are red?
The multiplication rule of probability of two dependent events is the product of the probability of first independent marginal event and the conditional probability of second dependent marginal event. Same applies to the joint probability of multiple dependent events.
Symbolically, let A be an independent event and B be a dependent event, the joint probability that both events and A and B occur is stated as:
P(A and B) = P(A)*P (B/A), where P(A) is the probability of an independent event A and P(B/A) is the conditional probability of a dependent event B given the occurrence of an independent event A.
The events are said to be dependent events if the occurrence of the first event affects the probability of the second and following events. In this example, drawing of a red card in the first event and not replacing back that card to the deck of cards before second draw of a card affects the second draw of a card. This applies to multiple draws of cards without replacement from a deck of cards.
Let R1 be an independent event of drawing a red card in the first draw of a card.  There are 26 red cards in a deck of cards. The marginal probability of R1, P(R1) is 26 divided by 52, 0.5. Let R2 be a dependent event of drawing a red card from the deck of cards in which the first drawn card is not replaced back.  There are 25 red cards left in the deck of 51 cards, as the first drawn cards is red which is not replaced back in the deck. Thus, the conditional probability of R2 given R1 in the first draw, P(R2/R1) is 25 divided by 51, 0.4902.
Now, applying the multiplication rule of dependent events, the joint probability that red cards are drawn in both draws without replacement is the product of P(R1) in the first draw and the P(R2) in the second draw.
Symbolically, P(R1 and R2) = P(R1)*P (R2/R1)=26/52 *25/51 =0.2450
The total number of outcomes in two draws of cards without replacement from a deck of cards and their probabilities are shown using a tree diagram below:









Figure 1: Marginal probabilities of first draws and conditional probabilities of second draws given the first draws of cards without replacement and joint probabilities of possible outcomes

Saturday, May 11, 2019

Time to Break Gender Stereotype of All Kinds

There could be gender stereotypes in some professions in the past leading people to perceive a man is a doctor and a woman a nurse, a man is a pilot and a woman an airhostess. However, there is nothing that men can do and women cannot or vice versa in this millennium. The constitution clearly prohibits against discriminating on the grounds of gender in remuneration and social security for the same work. A male doctor can be a gynecologist and a female doctor can treat a male reproductive organ.  Equal opportunity for men and women needs to be given in every profession. Thus, gender stereotypes need to be discouraged by creating awareness, creating opportunities and taking legal actions against the violation of the rules.
Basan Shrestha, Ghattekulo, Kathmandu

The Himalayan Times, PeopleSpeak, May 12, 2019
http://epaper.thehimalayantimes.com/html5/reader/production/default.aspx?pubname=&pubid=cd7278e2-4150-475f-8abe-305e5ed57783

Sunday, May 5, 2019

Multiplication Rule of Independent Events, Statistical Note 38

If an unbiased coin with two sides (Head and Tail) is tossed twice, what is the joint probability that head appears in both tosses?
The multiplication rule of probability of two independent events is the product of the probability of first independent marginal event and the probability of second independent marginal event. Same applies to the joint probability of multiple independent events.
Symbolically, if two events A and B are independent, the probability that both events and A and B occur is stated as:
P(A and B) = P(A)*P (B)

The events are said to be independent events if the occurrence of the first event does not affect the probability of the second and following events. In this example, occurrence of head or tail in the first toss does not affect the occurrence of head or tail in the second toss. Any of head or tail could appear in any number of toss of a coin. This applies to multiple tosses of a coin.

Let H be an independent event of turning head in a toss of a coin.  Likewise, let T be an independent event of turning tail in a toss of a coin.  The marginal probability of H, P(H) is one by two, 0.5. Similarly, the marginal probability of T, P(T) is one by two, 0.5.

The total number of outcomes in two tosses of a coin can be shown a tree diagram like the one shown below:










Figure 1: Marginal probabilities of independent events in first and second tosses of a coin and number of outcomes
Now, applying the multiplication rule of independent events, the joint probability that head appears in both tosses is the product of P(H) in the first toss and the P(H) in the second toss.
Symbolically, P(H and H) = ½ * ½ = ¼ = 0.25
Alternatively, the occurrence of head in both tosses is Outcome 1 of four outcomes in the tree diagram, and thus the probability is one out of four, equal to 0.25.

Sunday, April 28, 2019

Food of the People: Let's Eat!


I would introduce Buff Chilli to my foreign friends as a representative Nepali food. It is becoming popular in parties, restaurants and invitations in homes. However, I chose this dish based on my family’s experience. Previously, dishes prepared from buffalo meat or buff were consumed by Newars and several other ethnic groups, among others. But, in current days these dishes are increasingly popular among urban people, irrespective of any caste and ethnicity. Buff chili is prepared by first boiling pieces of buffalo meat and then chopping into pieces. Spices and other ingredients such as garlic, ginger, turmeric, cumin seed powder, chili powder and salt are added to the chopped meat and all stuffs are well mixed. The mixture of meat and spices is fried with onion and capsicum to soften buff and mix well with all spices and vegetables. Afterwards, tomato sauce is poured over the mixture before serving the dish.

Basan Shrestha, Ghattekulo, Kathmandu

Monday, October 8, 2018

What Sample Proportion is a Statistically Significant Estimate of Population Proportion?, Statistical Note 37

Toss a coin 20 times in a sample and count the number of heads. Repeat the same process for seven samples each constituting 20 tosses. Calculate the sample proportion of heads for each sample and test whether each sample proportion is a significant estimate of the population proportion. Discuss what makes the sample proportion a statistically significant estimate of the population proportion.

Key Words: Population Proportion, Sample Proportion, Standard Error, z-score, Statistical Significance, Sample Size

Introduction

Not all sample proportions are statistically significant estimates of the population proportion. Then, questions arise what sample proportion is significant and what makes the statistical significance. Tossing of a coin is an example of the binary categorical random variable to explain the statistical significance of the sample proportions. Refer to my statistical note 36 to know more about population and sample proportions and normality of the sampling distribution of the sample proportions.

Observed Data

I tossed a coin 20 times in a sample (S) and the same process was repeated for seven samples (S1 to S7). Can one guess how many heads will there be in each sample? Table 1 presents the outcome of 20 tosses of a coin in each of seven samples.

Table 1: Outcomes in 20 tosses of a coin in each of seven samples
















Every sample of 20 tosses in Table 1 is from a population constituting the large number of possible tosses. Below section discusses on whether the sample proportions are statistically significant estimates of the population proportion.

Discussion

Population proportion, denoted by ‘p’ in coin tossing experiments is 0.5. Sample proportion is denoted by  ’p^’ and spelt as ‘p-Hat’. The sample proportion of heads in 20 tosses of a coin ranged between 0.35 to 0.65 (Table 1).

Mean of sample proportions ‘p^’, also called center, is the population proportion ‘p’. Symbolically, it is indicated by µp^=p. In coin tossing, 0.5 is the mean of sample proportions or the population proportion.

Standard Deviation of Sample Proportions is expressed as the square root of the population proportion multiplied by one minus population proportion divided by sample size. This is referred to as spread or Standard Error (SE) of sample proportions, denoted by σp^ is Ö[p x (1-p)/n] where ‘p’ is the population proportion and ‘n’ is the sample size. In this example, SE is calculated to be Square root [0.5 x (1-0.5)/20], equal to 0.111803.

Sampling distribution of sample proportions with the sample size of 20 tosses of a coin is approximate to normal distribution with mean p=0.5 and σp^=0.111803.

In sampling distribution of sample proportions following normal distribution, z-score or test statistic is a measure calculated as the difference between population proportion and sample proportion divided by SE of sample proportions. Symbolically, z=(p+p^)/Ö[p x (1-p)/n]. In normal distribution the z- score equal to minus or plus 1.96 is a commonly used cut-off point for the sample proportion to be a statistically significant estimate of the population proportion indicating that 95 percent samples have population proportion within the confidence interval of minus 1.96 to plus 1.96. It means one is 95 percent confident that the population proportion will fall within 1.96 confidence interval. z-score ranging between negative to positive cut-off point is called the confidence interval. z-score less than minus 1.96 and greater than plus 1.96 indicate that the sample proportion is a statistically significant estimate of the population proportion from a different population. One point to note here is that the sample size ‘n’ is directly proportional to z-score indicating that as the sample size ‘n’ increases, z-score also increases.

I calculated the z-score for all seven sample proportions in this example (Table 2). For example, for the sample proportion P^=0.35, z = (p^-p) / σp^, where σp^=Ö[p x (1-p)/n]  = Ö[0.50 x 0.50/20] = 0.111803 so that z = (0.35-0.50) / 0.111803 = 1.3416. Likewise, the z-score was calculated for each sample proportion and tabulated. Using the cut-off point of z-score equal to minus or plus 1.96, the sample proportions in this example were not found to be the significant estimates of the population parameter of 0.50, given the sample size of 20 tosses of a coin.

Table 2: Sample proportions and their significance to estimate population proportion of samples constituting 20 tosses of a coin









The sample proportions as smaller as 0.2808 and as bigger as 0.7192 are statistically significant estimates of the population proportion 0.50, given the sample size of 20 tosses of a coin.

If the sample size is increased, the sample proportions bigger than 0.2808 and smaller than 0.7192 will be statistically significant estimates of population proportion of 0.50. Example, if the sample size is increased to 50 tosses, the sample proportions as smaller as 0.3614 and as bigger as 0.6386 are statistically significant estimates of the population proportion of 0.50. Likewise, if the sample size is increased to 100 tosses, the sample proportions as smaller as 0.4020 and as bigger as 0.5980 are statistically significant estimate of the population proportion of 0.50.

I am curious whether John Kerrich’s observed sample proportion of heads equal to 0.5067 in 10,000 tosses of a coin is a statistically significant estimate of population proportion of 0.5 or not. z-score for this is calculated to be 1.34, which is lower than the cut-off point of 1.96 indicating that this is one of 95 percent samples each of 10,000 tosses so that one can be 95 percent confident that this sample proportion is an insignificant estimate of the population proportion, 0.50. Thus, this sample proportion is not a statistically significant estimate of the population proportion, 0.50.

Conclusion

Not all sample proportions are statistically significant estimates of the population proportion. The sample size is pivotal for identifying whether the sample proportion is a statistically significant estimate of the population proportion or not.